40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2602     Entry time: Mon Jul 12 14:42:42 2021
Author: Stephen 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: RTD attached to coldhead with spring clamp, Si mass to be installed this week 

Brief summaries of the last week's progress and the coming week's plans (plots will be posted soon!):

- progress Friday 09 July: Opened the cryostat up at the cold head, and attached an RTD to the cold head with a spring clamp (instead of relying purely on the cryo varnish).

- progress Monday 12 July: Found 65 K workpiece temp and 63 K cold head temp. RTD was apparently held successfully by spring clamp, and we will continue to collect cold head temperature in future runs. Warmup was started, with old data collection completed (cooldown_20210709) and new data collection commenced (warmup_2021_07_12). Note that warmup started at 1:14 pm, and it took me ~ 5 minutes to stop and restart the script to changeover to the warmup data collection.

- table plan Wednesday 15 July: Complete in-air optical layout. Make one flat face of Si mass reflective.

- chamber plan Thursday 16 July: Open up main volume and drop in frame with Si mass. Connect RTDs. Start cooldown. Confirm cooldown is going ok (optical alignment, especially), and revert if necessary before things get too cold.

- table plan Friday 16 July: Maybe measure stuff, maybe better to wait till coming week and use controlled heating to hit different temperature setpoints.

 

ELOG V3.1.3-